OpenStudy (anonymous):

Bill keeps a colony of honeybees on his property. He has created an expression to represent the number of honeybees that are in the colony t years after he decided to create the colony. The expression Bill created is -16t4 + 160t2 – 144. After deciding to create the colony, it took a fair amount of time to prepare before actually getting bees in his colony. After some time, the population in the colony started to decrease; eventually Bill lost his colony. a. How could you use Bill’s expression to find out when Bill first got his honeybees and when he lost his colony?

OpenStudy (anonymous):

Use the caret ^ (above the 6) to denote an exponent !! Bill's expression: [- 16t^4 + 160t^2 - 144] set Bill's expression for the colony equal to ZERO to find the time (t) when there are zero bees... - 16t^4 + 160t^2 - 144 = 0 - 16(t^4 - 10t^2 + 9) = 0 solve the equation (in quadratic form)... (t^2 - 9)(t^2 - 1) = 0 t = ± 3 [- 3 is extraneous] or t = ± 1 [- 1 is extraneous] Bill got his colony after 1 year...Bill lost his colony after 3 years

OpenStudy (anonymous):

b. Factor out a negative 1 from the expression using a bracket. Then rewrite the first term inside the bracket as the square of a term. Do this so that if the term being squared were replaced with a variable, the term would have no coefficient.

OpenStudy (anonymous):

@Icedragon50

OpenStudy (anonymous):

._.

OpenStudy (anonymous):

i need to finish my exam srry maybe @Jaynator495 @JayJayTheWereWolfGal1 can help. (U BETTER HELP HULK)

OpenStudy (anonymous):

@icedragon50, nice plagiarism,. You could of at least gave credit to the person who did the work. https://answers.yahoo.com/question/index?qid=20150518060309AA1WMd0

OpenStudy (anonymous):

Please help... Factor out a negative 1 from the expression using a bracket. Then rewrite the first term inside the bracket as the square of a term. Do this so that if the term being squared were replaced with a variable, the term would have no coefficient.

OpenStudy (anonymous):

Did you ever get the answer?